OpenStudy (anonymous):

Sriram it says if the same wave has a frequency of 10Hz what is its wavelentgh?

OpenStudy (sriram):

look first you get the velocity in the first part then use vel.=wavelength*frequency

OpenStudy (anonymous):

so i use one of those??? for example wavelenth-frequency

OpenStudy (sriram):

i didn't get u...

OpenStudy (turingtest):

You have not provided enough info yet. Is it by any chance a light wave? we need the velocity of the wave to answer.

OpenStudy (turingtest):

though 10Hz is pretty low for light lol

OpenStudy (anonymous):

okay maybe it says the spped formula is wave length frequency or speed and it says you can calculate the third??

OpenStudy (anonymous):

then it asks that

OpenStudy (turingtest):

the what formula? I think we are all confused by your exact question

OpenStudy (anonymous):

the speed? here i did the expermant il upload the file hold on.. ingore the expermanet

OpenStudy (anonymous):

????

OpenStudy (turingtest):

ok so the equation is \[v=\lambda f\]where v=velocity lambda=wavelength, (the greek letter) f=frequency now we get do make up a wavelength... Go ahead, make one up!

OpenStudy (sriram):

damn it it says 6meters in 2 seconds not 6m & 2s

OpenStudy (anonymous):

Can you fix it lol???

OpenStudy (turingtest):

that's okay, we get to make up our own wavelength, so v=6/2=3m/s and lambda is whatever we ant it to be, so we can solve it

OpenStudy (turingtest):

want*

OpenStudy (anonymous):

it will be wavelentgh?

OpenStudy (turingtest):

yes give me a number!

OpenStudy (turingtest):

make it up!

OpenStudy (anonymous):

8

OpenStudy (turingtest):

ok then 8m lambda=8 v=3 solve for f!

OpenStudy (sriram):

speed=wavelength* frequency 3=wavelength*10 hence wavelength=3/10=0.3meters

OpenStudy (anonymous):

so its that?

OpenStudy (turingtest):

oh I see it asks for the definition of wavelength, not to define one :( nevermind I misread

OpenStudy (anonymous):

can you fill out the sheet for me and send it to me back?? lol

OpenStudy (turingtest):

nope gotta learn buddy :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

tell me the ansnwers lol..

OpenStudy (turingtest):

not what I'm here for, look for someone else

OpenStudy (anonymous):

wow... Siriam?